OpenStudy (riley):

Gymnast Clothing manufactures expensive soccer cleats for sale to college bookstores in runs of up to 500. Its cost (in dollars) for a run of x pairs of cleats is: C(x) = 2950 + 6x + 0.1x2 (0 ≤ x ≤ 500). Gymnast Clothing sells the cleats at $130 per pair. Find the revenue and profit functions. How many should Gymnast Clothing manufacture to make a profit? I've worked this problem out so many times I'm almost sick of looking at it, and I still haven't come up with the correct answer.

OpenStudy (riley):

This is what I keep coming up with: Revenue: R(x) = 130x Profit: P(x) = .2950+124x-0.1x^2 And for the pairs of shoes I'm getting 23.5. When I was working through the profit formula however, to find the break-even point I set the equation equal to 0, and then solved for x. I ended up with a quadratic equation so I used the quadratic formula to solve that part. I ended up getting: = 23.5 or -1263.5; And because of the domain being from 0 to 500 I assumed the answer had to be 23.5 . . . however this isn't correct and I can't figure out where I'm messing up at. Could someone please help.

OpenStudy (riley):

-2950+124x-0.1x^2 typo**

OpenStudy (riley):

I still need help, please someone. I just worked through this problem for the fifth time, and I'm stuck. I can't figure out where I'm going wrong at and this is my last problem to finish. >_< uughh

OpenStudy (gw2011):

R=130x P=R-C P=130x-(2950+6x+0.1x^2 P=124x-2950-0.1x^2 P=-0.1x^2+124x-2950 I used the quadratic formula to solve for x: x=[-b+/-sqrt(b^2-4ac)]/2a a=-0.1 b=124 c=-2950 x=[-124+/-sqrt(124^2-(4)(-0.1)(-2950)]/(2)(-0.1) x=[-124+/-sqrt(15,376-1180)]/-0.2 x=[-124+/-sqrt(14,196)]/-0.2 x=[(-124)+/-(119.1)]/-0.2=[(-124)+/-(119)]/-0.2 x=(-124+119)/-0.2=-5/-0.2=25 x=(-124-119)/-0.2=-243/-0.2=1215 which is not a solution because 0<=x<=500 Therefore, 25 pairs of cleats need to be made to make a profit

OpenStudy (riley):

Wow, I was doing the problem the exact same way, however I see where I messed up at. When I was using the quadratic formula to solve for x, I didn't carry over one of my signs correctly, and it screwed everything up. Thank you so much for showing me that, I can't believe one little minor mistake like that caused so much trouble. Thank you again!!<3

OpenStudy (gw2011):

You're welcome. I'm glad I could help.